# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v1

@@ 5.3 Comparisons with prior work
Table 4: Operator perspective results. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the number of trips satisfied with number of transfers $i$, while $d_{un}$ is the number of trips satsified with 3 or more transfers. Arrows next to each quantity indicate which of increase or decrease is desirable. Bolded values in $C_{o}$, $d_{0}$ and $d_{un}$ columns are the best on that environment.
@@
The percentage-per-transfers metrics $d_{0}$ to $d_{un}$ reveal that both PC-EA and NEA favour higher numbers of transfers relative to most of the other methods, particularly Hüsselmann et al. (2023). This is true for both the passenger and operator perspectives. For the operator perspective, this matches our intuitions: shorter routes necessitate more transfers to achieve full coverage. But we were surprised to observe this in the passenger perspective case as well. Compared to Hüsselmann et al. (2023)’s DBMOSA, PC-EA achieves lower mean passenger travel time on Mumford2 and Mumford3 for the passenger perspective, despite the fact that it requires more passenger trips to make 1 and 2 transfers. Still, PC-EA does give fewer transfers overall than NEA, which matches our intuition that fewer transfers should result in lower average trip time $C_{p}$.
@@
It seems that the process of compositing shortest paths to form routes biases those routes towards directness and efficiency, to a degree that outweighs the cost imposed by having more transfers. Each transfer is less impactful on riders if the routes they’re transferring between are each more direct. Meanwhile, since both the learned heuristic of NEA and the unlearned heuristic of PC-EA suffer from high transfer counts, it seems the difference between these and the other results from the literature may be a result of the evolutionary algorithm itself. At any rate, these results are not out of line with those of Mumford (2013a) and Kılıç and Gök (2014) on the largest cities, so we do not consider this a major strike against our heuristics. Indeed, we expect that these results could be improved by explicitly including $d_{0}$ in the cost function used to train the neural net policy, and this would reduce the number of transfers achieved by NEA.
6 REAL-WORLD EVALUATION
6.1 Laval Scenario
@@
To assess our method in a realistic problem instance, we applied it to a dataset representing the city of Laval in Quebec, Canada. Laval is a suburb of the major Canadian city of Montreal. As of the 2021 census, it had a total population of 429,555 (Statistics Canada, 2023). Our model of the city of Laval is based on several sources: geographic data on census dissemination areas for 2021 from Statistics Canada (Statistics Canada, 2021), a map of the street network of the city, an OD dataset (Agence métropolitaine de transport, 2013), and publicly-available GTFS data from 2013 (STL, ) that describes the existing transit system in the city.
@@ 6.1.1 Street graph
The street graph ${\mathcal{N},\mathcal{E}_{s}}$ was derived from the 2021 census dissemination data by taking the centroid of each dissemination area within Laval to be a node in $\mathcal{N}$, and adding a street edge between any two nodes whose dissemination areas shared a border. To assign drive times $t_{ij}$, we find all points in the road network within the dissemination areas of $i$ and $j$, compute the shortest-path driving distance over the road network between each pair of points in $i$ and $j$, and set $t^{\prime}_{ij}$ as the median of these driving distances. We do this so that its value will better reflect the real drive times given the existing road network. As our method treats cities as undirected graphs, it expects that $t_{ij}=t_{ji}\;\forall\;i,j\in\mathcal{N}$. To enforce this, as a final step we set:
@@ 6.1.2 Existing transit
The existing transit system in Laval has 43 bus routes that operate during morning rush hour from 7 to 9 AM. We map each stop on an existing bus route to a node in $\mathcal{N}$ based on which census dissemination area contains the stop. Doing this for all routes gives us a version of the existing transit system that runs over our street graph ${\mathcal{N},\mathcal{E}_{s}}$, allowing us to compare it directly with solutions output by our algorithm. We will refer to this mapped transit system as the STL network, after the city’s transit agency “Société de transport de Laval”.
The numbers of stops on the routes in the STL network range from 2 to 52. To ensure a fair comparison, we set $MIN=2$ and $MAX=52$ in the same way when running our algorithm. Unlike the routes in our algorithm’s solutions, many of STL’s routes are not symmetrical (that is, they follow a different path in each direction between terminals), and several are unidirectional (they go only one way between terminals). We maintain the constraints of symmetry and bidirectionality on our own algorithm, but do not enforce them upon the STL network. The only modification we make is that when reporting total route time $C_{o}$ in this section, we calculate it by summing the travel times of both directions of each route, instead of just one direction as in the preceding sections.
In addition, there is an underground metro line, the Montreal Orange Line, which has three stops in Laval. As with the exisiting bus lines, we map these three stops to their containing DAs and treat them as an additional bus route. We consider it to be fixed, and we include it in our evaluation both of the STL network and the solutions produced by our algorithm.
@@ 6.1.3 Demand matrix
Each entry in the OD dataset corresponds to a trip, and has (lat,lon) coordinates $l_{o}$ and $l_{d}$ for the trip’s origin and destination, as well as an “expansion factor” $e$ giving the estimated number of actual trips corresponding to that surveyed trip. It also indicates the mode of travel, such as car, bicycle, or public transit, that was used to make the trip.
To assemble $D$ for Laval, we initialize its entries to 0. Then for each entry in the OD dataset, we find which dissemination area contains each of $l_{o}$ and $l_{d}$, and associate them with the matching nodes $i$ and $j$ in $\mathcal{N}$. We then update $D_{ij}\leftarrow D_{ij}+e$, so that $D$ has the estimated demand between every pair of census dissemination areas. To enforce symmetric demand, we then assign $D_{ij}\leftarrow\max(D_{ij},D_{ji})\;\forall\;i,j\in\mathcal{N}$.
@@
Many entries in the OD dataset refer to trips that begin or end in Montreal. For our purposes, we “redirect” all trips that enter or leave Laval and are made by public transit to one of several “crossover points” that we define. These crossover points are the locations of the three Orange Line metro stations in Laval, and the locations of the last or first stop in Laval of each existing bus line that goes between Laval and Montreal. For each trip between Laval and Montreal, we identify the crossover point that has the shortest distance to either of the trip’s endpoints, and replace the Montreal end-point of the trip by this crossover point. This process was automated with a simple computer script. Our intuition here is that if the Laval end of the trip is close to a transit stop that provides access to Montreal, the rider will choose to cross over to the Montreal transit system as soon as possible, as most locations in Montreal will be easier to access once in the Montreal transit system. Trips that go between Laval and Montreal by means other than public transit are not included in $D$. The resulting demand matrix $D$ contains 548,159 trips, of which 63,104 are trips between Laval and Montreal that we redirected.
@@
Having done the above, we apply one last filtering step. The STL network does not provide a path between all node-pairs with non-zero demand in the demand matrix $D$, and so violates constraint 1 of section 3. This is because $D$ at this point includes all trips that were made by any means, including to and from areas that are not served by the STL network, possibly because they are populated by car owners unlikely to use transit if it were available. To ensure a fair comparison between the STL network and our algorithm’s solutions, we set to 0 all entries of $D$ for which the STL network does not provide a path. By doing this, we also cause our system to output transit networks that are “closer” to the existing transit network, in that they satisfy the same travel demand as before. This should reduce the scope of the changes required to go from the existing network to a new one proposed by our system.
This way, we consider only demand that is likely to use transit when it is available, because it corresponds to residents already living close to transit. Table 5 contains the statistics of, and parameters used for, the Laval scenario.
@@
[…]
@@
Figure 6: At left, a map of the census dissemination areas of the city of Laval, with “crossover points” used to remap external demands shown as black circles, as well as the portion of the Orange Line metro that we include. At right, the street graph constructed from these dissemination areas.
@@ 6.2 Enforcing constraint satisfaction
In the experiments on the Mandl and Mumford cities described in section 5, the NEA algorithm consistently produced solutions that did not violate any of the constraints in section 3. However, the Laval scenario is considerably larger than even the largest Mumford scenario (632 nodes versus 127), and in our initial experiments we found that both EA and NEA gave solutions that violated constraint 1, providing no transit path for some node pairs $(i,j)$ for which $D_{ij}>0$. That is, constraint 1 requires that $c_{1}(\mathcal{R})=0$, where $c_{1}(\mathcal{R})$ is the number of unconnected node-pairs with non-zero demand over transit network $\mathcal{R}$.
To remedy this, we modified the MDP described in subsection 3.1 to only allow actions that reduce $c_{1}$, if any such action is available. Specifically, the following changes are applied to the action space $\mathcal{A}_{t}$ whenever $c_{1}(\mathcal{R}_{t})>0$:
@@
  • When $t$ is odd, if $\mathcal{A}_{t}$ contains any paths that would reduce $c_{1}(\mathcal{R}_{t})$ if added to $r_{t}$, paths that would not reduce $c_{1}(\mathcal{R}_{t})$ are excluded from $\mathcal{A}_{t}$. This means that if it is possible to connect some unconnected demand by extending a route, then not doing so is disallowed.
@@
We apply the same changes when the neural mutator in NEA and its variants plans an individual route. We find that by using these changes both when generating the initial solution via LC-100 and during the subsequent NEA run, the resulting solutions have no constraint violations, even when using policy weights $\theta$ that were trained without these changes to the MDP. The results reported in this section are obtained using this variant of the MDP.
6.3 Results
$\alpha$ Method $C_{p}$ $C_{o}$ $d_{0}$ $d_{1}$ $d_{2}$ $d_{un}$
STL 124.61 23954 14.48 22.8 20.88 41.83
LC-100 88.83 21893 14.19 23.45 23.58 38.78
0.0 NEA 96.65 19624 14.22 23.57 23.63 38.57
PC-EA 67.67 31907 16.56 32.94 29.06 21.44
LC-100 79.32 24124 14.88 24.87 24.53 35.72
0.5 NEA 82.55 20978 14.40 24.44 24.76 36.41
PC-EA 67.40 31784 16.33 33.35 28.99 21.33
LC-100 66.49 32270 16.35 29.77 26.04 27.85
1.0 NEA 59.23 45382 20.99 34.48 26.71 17.82
PC-EA 57.42 49016 22.89 39.39 26.33 11.38
Table 6: Performance of STL network, and our system’s networks at three $\alpha$ values. Bolded values are best for a given $\alpha$. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the number of trips satisfied with number of transfers $i$, while $d_{un}$ is the number of trips satsified with 3 or more transfers. All values are averaged over ten random seeds, excluding STL which is a fixed network.
@@ 6.3 Results
Some trial runs with $IT=4,000$ showed that the solution quality stopped improving after $2,700$ iterations, so to save time we ran our final NEA and PC-EA experiments on Laval with $IT=3,000$. Other parameters are the same as in the Mumford experiments: $B=10$, $E=10$, and transfer penalty $p_{T}=300$s. We perform three sets of experiments from three perspectives: operator ($\alpha=0.0$), passenger ($\alpha=1.0$), and balanced ($\alpha=0.5$) As above, we perform ten runs of each experiment with ten different random seeds and ten different sets of trained policy weights; the same per-seed weights are used as in the previous experiments.
Figure 7: Trade-offs between average trip time $C_{p}$ and total route time $C_{o}$ achieved by LC-100, NEA, PC-EA, for $\alpha=0.0$, 0.5, and 1.0, as well as the trade-off for the STL network itself. Each point is a mean value over 10 random seeds, and error bars show one standard deviation. Note that PC-EA’s $\alpha=0.0$ and 0.5 have nearly identical values (as shown in Table 6), so the corresponding points are stacked in this figure and appear as one point.
@@
The results of all experiments are shown in Table 6. As the STL’s network is fixed, its metrics are independent of $\alpha$, so we show it only once. The results for the LC-100 solution that are used to initialize NEA are presented as well, to show how much improvement the NEA makes over its initial solution. Figure 7 highlights the trade-offs each method achieves between average trip time $C_{p}$ and total route time $C_{o}$, and how they compare to the STL network.
@@
Comparing NEA’s results to LC-100, we see that at $\alpha=0.0$, NEA decreases $C_{o}$ and increases $C_{p}$ relative to LC-100; and for $\alpha=1.0$, the reverse is true. In both cases, NEA improves over the initial solution on the sole objective being optimized, at the cost of the other, as we would expect. In the balanced case ($\alpha=0.5$), we see that NEA decreases $C_{o}$ by 13% while increasing $C_{p}$ by 4%. This matches what we observed in Figure 3, as discussed in subsection 5.1.
@@
Looking at the transfer percentages $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$, we see that the overall number of transfers shrinks, with $d_{1}$ growing and $d_{un}$ shrinking, as $\alpha$ increases. This is as expected, since fewer transfers usually means shorter passenger travel times. By the same token, PC-EA performs best on these metrics, as it does on average trip time $C_{p}$, since the two are causally linked. Notably though, at $\alpha=0.0$ NEA’s $d_{0}$, $d_{1}$, and $d_{2}$ are only slightly higher than the STL network’s - the largest increase is in $d_{2}$, by 13% of STL’s value - while $d_{un}$ is less than the STL network’s, meaning that NEA’s output networks let more passengers reach their destination in 2 or fewer transfers than the STL network. In fact, given the very small decrease in $d_{0}$, the increase in $d_{1}$ and $d_{2}$ mostly comes from passengers who contributed to $d_{un}$ in the STL network; NEA has thus decreased the average number of transfers. This is a significant improvement, given that at $\alpha=0.0$ the algorithm is unconcerned with minimizing passenger travel time.
@@
At $\alpha=0.5$, NEA’s $d_{un}$ decreases further and $d_{0}$, $d_{1}$, and $d_{2}$ correspondingly increase, though the change in these values from $\alpha=0.0$ is very small relative to the decrease in $C_{p}$. Then from $\alpha=0.5$ to $1.0$, NEA’s $d_{un}$ drops by more than half, and its $d_{0}$, $d_{1}$, and $d_{2}$ are all markedly increased, especially in $d_{0}$ which is 31% greater than STL’s $d_{0}$.
The likely aim of Laval’s transit agency is to reduce costs while improving service quality, or at least minimizing its decline. The “balanced” case, where $\alpha=0.5$, is most relevant to that aim. In this balanced case, we find that NEA proposes solutions that reduce total route time by 13% versus the existing STL transit network.
@@
Some simple calculations can help understand the savings this could represent. The approximate cost of operating one of Laval’s buses is 200 Canadian dollars (CAD) per hour. Suppose that each route has the same headway (time between bus arrivals) $H$. Then, the number of buses $N_{r}$ required by route $r$ is the route’s driving time $\tau_{r}$ divided by the headway, $\frac{\tau_{r}}{H}$. The total number of buses $N_{\mathcal{R}}$ required for a transit network $\mathcal{R}$ is then:
$N_{\mathcal{R}}=\sum_{r\in\mathcal{R}}N_{r}=\frac{1}{H}\sum_{r\in\mathcal{R}}\tau_{r}=\frac{C_{o}}{H}$
@@
So the total cost of the system is 200 CAD×$\times N_{\mathcal{R}}$. Assuming a headway of 15 minutes on all routes, that gives a per-hour operating cost of 319,387 CAD for the STL network, while the average operating cost of networks from NEA with $\alpha=0.5$ is 279,707 CAD. So under these simplifying assumptions, NEA could save the Laval transit agency on the order of 40,000 CAD per hour - savings of more than 12% of costs. This is only a rough estimate: in practice, headways differ between routes, and there may be additional costs involved in altering the routes from the current system, such as from moving or building bus stops. But it shows that NEA, and neural heuristics more broadly, may offer practical savings to transit agencies.

### v2

@@ 5.3.2 Operator-perspective results
Table 4: Operator perspective results. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satsified with 3 or more transfers. Arrows next to each quantity indicate which of increase or decrease is desirable. Bolded values in $C_{o}$, $d_{0}$ and $d_{un}$ columns are the best on that environment.
@@
The metrics $d_{0}$ to $d_{un}$ reveal that both PC-EA and NEA favour higher numbers of transfers relative to most of the other methods, particularly Hüsselmann et al. (2023). This is true for both the passenger and operator perspectives. For the operator perspective, this matches our intuitions: shorter routes will deliver fewer passengers directly to their destinations. But we were surprised to observe this in the passenger perspective case as well.
Compared to Hüsselmann et al. (2023)’s DBMOSA, PC-EA achieves lower average trip time $C_{p}$ on Mumford2 and Mumford3 for the passenger perspective, despite the fact that it requires more passenger trips to make 1 and 2 transfers. Still, PC-EA does give fewer transfers overall than NEA, which matches our intuition that fewer transfers should result in lower average trip time $C_{p}$. It seems that the process of compositing shortest paths to form routes biases those routes towards directness and efficiency, to a degree that outweighs the cost imposed by having more transfers. Each transfer is less impactful on riders if the routes they’re transferring between are more direct.
@@
Meanwhile, since both the learned heuristic of NEA and the unlearned heuristic of PC-EA suffer from high transfer counts, it seems the difference between these and the other results from the literature may be a result of the evolutionary algorithm itself. At any rate, these results are not out of line with those of Mumford (2013a) and Kılıç and Gök (2014) on the largest cities, so we do not consider this a major strike against our learned heuristics. Indeed, we expect that these results could be improved by explicitly including $d_{0}$ in the reward function used to train the policy $\pi_{\theta}$, and this would reduce the number of transfers required by NEA’s networks.
6 LAVAL EXPERIMENTS
@@
To assess our method in a realistic problem instance, we applied it to data from the city of Laval in Quebec, Canada. Laval is a suburb of the major Canadian city of Montreal. As of the 2021 census, it had a total population of 429,555 (Statistics Canada, 2023). Public transit in Laval is provided primarily by the bus network of the Société de Transport de Laval; additionally, one line of the Montreal underground metro system has three stops in Laval.
6.1 Representing Laval
To apply our learned policy $\pi_{\theta}$ to Laval, we first had to assemble a realistic city graph $\mathcal{C}=(\mathcal{N},\mathcal{E}_{s},D)$ representing Laval from the data sources that were available to us. Our model of the city of Laval is based on several sources: geographic data on census dissemination areas for 2021 from Statistics Canada (Statistics Canada, 2021), a GIS representation of the road network of Laval provided to us by the Société de Transport de Laval, an OD dataset (Agence métropolitaine de transport, 2013) provided by Quebec’s Agence Métropolitaine de Transport, and publicly-available GTFS data from 2013 (STL, ) that describes Laval’s existing transit system.
@@ 6.1.1 Street graph
The street graph $(\mathcal{N},\mathcal{E}_{s})$ was derived from the 2021 census data by taking the centroid of each CDA within Laval to be a node in $\mathcal{N}$, and adding a street edge $(i,j,t_{ij})$ for node pair $(i,j)$ if their corresponding CDAs shared a border. To compute drive times $t_{ij}$, we found all points in the road network within the CDAs of $i$ and $j$, computed the shortest-path driving time over the road network from each of $i$’s road-network points to each of $j$’s, and set $t^{\prime}_{ij}$ as the median of these driving times. We did this so that $t^{\prime}_{ij}$ would reflect the real drive times given the existing road network. By this approach, $t^{\prime}_{ij}\neq t^{\prime}_{ji}$ in general, but as our method treats cities as undirected graphs, it expects that $t_{ij}=t_{ji}\;\forall\;i,j\in\mathcal{N}$. To enforce this, as a final step we set:
@@ 6.1.2 Existing transit
The existing transit network in Laval has 43 bus routes that operate during morning rush hour from 7 to 9 AM. To compare networks from our algorithm to the this network, we had to translate it to a network that runs over $(\mathcal{N},\mathcal{E}_{s})$ as defined in subsection 6.1.1. To do this, we mapped each stop on an existing bus route to a node in $\mathcal{N}$ based on which CDA contains the stop. We will refer to this translated transit system as the STL network, after the city’s transit agency, the Société de Transport de Laval.
The numbers of stops on the routes in the STL network range from 2 to 52. To ensure a fair comparison, we set $MIN=2$ and $MAX=52$ when running our algorithm. Unlike the routes in our algorithm’s networks, many of STL’s routes are not symmetric (that is, they follow a different path in each direction between terminals), and several are unidirectional (they go only one way between terminals). We maintained the constraints of symmetry and bidirectionality on our own algorithm, but did not enforce them upon the STL network. The only modification we made is that when reporting total route time $C_{o}$ in this section, we calculated it for bidirectional routes by summing the travel times of both directions, instead of just one direction as in the preceding sections.
In addition, there is an underground metro line, the Montreal Orange Line, which has three stops in Laval. As with the existing bus lines, we mapped these three stops to their containing CDAs and treat them as forming an additional route. We added this metro route to both the STL network and the networks $\mathcal{R}$ produced by our algorithm when evaluating them, to reflect that unlike the bus routes, it is not feasible to change the metro line and so it represents a constant of the city.
@@ 6.1.3 Demand matrix
Finally we assembled the demand matrix $D$ from the OD dataset. The entries in the OD dataset correspond to trips reported by residents of Laval who were surveyed about their recent travel behaviour. Each entry has (lat,lon) coordinates $l_{o}$ and $l_{d}$ for the trip’s approximate origin and destination, as well as an “expansion factor” $e$ giving the estimated number of actual trips corresponding to that surveyed trip. It also indicates the mode of travel, such as car, bicycle, or public transit, that was used to make the trip.
@@
Many entries in the OD dataset refer to trips that begin or end in Montreal. For our purposes, we “redirected” all trips made by public transit that enter or leave Laval to one of several “crossover points” that we defined. These crossover points are the locations of the three Orange Line metro stations in Laval, and the locations of the last or first stop in Laval of each existing bus route that goes between Laval and Montreal.
For each trip between Laval and Montreal, we identified the crossover point that has the shortest distance to either of $l_{o}$ or $l_{d}$, and overwrote the Montreal end-point of the trip by this crossover point. This process was automated with a simple computer script. The idea is that if the Laval end of the trip is close to a transit stop that provides access to Montreal, the rider will choose to cross over to the Montreal transit system as soon as possible, as most locations in Montreal will be easier to access once in the Montreal transit system. Trips that go between Laval and Montreal by means other than public transit were not included in $D$, as we judge that most such trips cannot be induced to switch modes to transit.
Having done this remapping, we initialized $D$ with all entries set to 0. Then for each entry in the OD dataset, we found the CDAs that contain $l_{o}$ and $l_{d}$, associated them with the matching nodes $i$ and $j$ in $\mathcal{N}$, and updated $D_{ij}\leftarrow D_{ij}+e$. $D$ then had the estimated demand between every pair of CDAs. To enforce symmetric demand, we then assigned $D_{ij}\leftarrow\max(D_{ij},D_{ji})\;\forall\;i,j\in\mathcal{N}$. The resulting demand matrix $D$ contains 548,159 trips, of which 63,104 are trips between Laval and Montreal that we redirected.
@@
We then applied a final filtering step. The STL network does not provide a path between all node-pairs $i,j$ for which $D_{ij}>0$, and so it violates constraint 1 of section 3. This is because $D$ at this point includes all trips that were made within Laval by any mode of transport, including to and from areas that are not served by the STL network. These areas may be unserved because they are populated by car owners unlikely to use transit if it were available. To ensure a fair comparison between the STL network and our algorithm’s networks, we set to 0 all entries of $D$ for which the STL network does not provide a path. We expected this to cause our system to output transit networks that are “closer” to the existing transit network, in that they satisfy the same travel demand as before. This should reduce the scope of the changes required to go from the existing network to a new one proposed by our system.
Table 5 contains the statistics of, and parameters used for, the Laval scenario.
@@
# nodes $n$ street edges $|\mathcal{E}_{s}|$ demand trips # routes $S$ $MIN$ $MAX$ Area (km2)
@@
Figure 6: 6(a) shows a map of the census dissemination areas of the city of Laval, with “crossover points” used to remap inter-city demands shown as black circles, and the Montreal Metro Orange Line shown in orange. 6(b) shows the street graph constructed from these dissemination areas.
@@ 6.2 Enforcing constraint satisfaction
In the experiments on the Mandl and Mumford benchmark cities described in section 5, the EA and NEA algorithms never produced networks violated any of the constraints on networks outlined in section 3. However, Laval’s city graph, with 632 nodes, is considerably larger than even the largest Mumford city, which has only 127 nodes. In our initial experiments on Laval, we found that both EA and NEA consistently produced networks that violated constraint 1, providing no transit path for some node pairs $(i,j)$ for which $D_{ij}>0$.
To remedy this, we modified the MDP described in subsection 3.1 to enforce reduction of unsatisfied demand. Let $c_{1}(\mathcal{C},\mathcal{R})$ be the number of node-pairs $i,j$ with $D_{ij}>0$ for which network $\mathcal{R}$ provides no transit path, and let $\mathcal{R}^{\prime}_{t}=\mathcal{R}_{t}\cup\{r_{t}\}$, the network formed from the finished routes and the in-progress route $r_{t}$. We applied the following changes to the action space $\mathcal{A}_{t}$ whenever $c_{1}(\mathcal{R}_{t})>0$:
@@
  • When $t$ is odd, if $\mathcal{A}_{t}$ contains any paths that would reduce $c_{1}(\mathcal{C},\mathcal{R}^{\prime}_{t})$ if added to $r_{t}$, then all paths that would not reduce $c_{1}(\mathcal{C},\mathcal{R}^{\prime}_{t})$ are removed from $\mathcal{A}_{t}$. This means that if it is possible to connect some unconnected trips by extending $r_{t}$, then not doing so is forbidden.
@@
With these changes to the MDP, we found that both LC-100 and NEA produced networks that connected all desired passenger trips, with no constraint violations - even when the policy $\pi_{\theta}$ was trained without these changes. The results reported in this section are obtained using this variant of the MDP.
6.3 Experiments
We performed experiments evaluating both the NEA and PC-EA algorithms on the Laval city graph. Three sets of experiments were performed with different $\alpha$ values representing different perspectives: the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). As in section 5, we perform ten runs of each experiment with ten different random seeds and ten different learned policies $\pi_{\theta}$, and report statistics over these ten runs; the same per-seed parameters $\theta$ are used as in the experiments of section 5. We also run the STL network on the Laval city graph $\mathcal{C}$ in the same way that we run the networks from our other algorithms to measure their performance, but since the STL network is a single, pre-determined network that does not depend on $\alpha$, we run it only once.
@@ 6.3 Experiments
Some trial runs we performed with $IT=4,000$ showed that the network cost stopped decreasing after $2,700$ iterations, so to save time we ran our final NEA and PC-EA experiments on Laval with $IT=3,000$. The other parameters of the algorithm were the same as in the Mumford experiments: $B=10$, $E=10$, and transfer penalty $p_{T}=300$s.
6.3.1 Results
@@
The results of the Laval experiments are shown in Table 6. The results for the initial networks produced by LC-100 are presented as well, to show how much improvement the NEA makes over its initial networks. Figure 7 highlights the trade-offs each method achieves between average trip time $C_{p}$ and total route time $C_{o}$, and how they compare to the STL network.
$\alpha$ Method $C_{p}\downarrow$ $C_{o}\downarrow$ $d_{0}\uparrow$ $d_{1}$ $d_{2}$ $d_{un}\downarrow$
N/A STL 124.61 23954 14.48 22.8 20.88 41.83
LC-100 88.83 $\pm$ 11.53 21893 $\pm$ 5096 14.19 $\pm$ 1.28 23.45 $\pm$ 3.97 23.58 $\pm$ 2.71 38.78 $\pm$ 7.52
0.0 NEA 96.65 $\pm$ 12.10 19624 $\pm$ 2455 14.22 $\pm$ 0.97 23.57 $\pm$ 3.67 23.63 $\pm$ 2.41 38.57 $\pm$ 6.39
PC-EA 67.67 $\pm$ 1.37 31907 $\pm$ 1295 16.56 $\pm$ 0.65 32.94 $\pm$ 1.09 29.06 $\pm$ 0.92 21.44 $\pm$ 2.04
LC-100 79.32 $\pm$ 8.27 24124 $\pm$ 7736 14.88 $\pm$ 1.98 24.87 $\pm$ 5.22 24.53 $\pm$ 1.85 35.72 $\pm$ 8.49
0.5 NEA 82.55 $\pm$ 7.11 20978 $\pm$ 3424 14.40 $\pm$ 1.35 24.44 $\pm$ 4.16 24.76 $\pm$ 2.17 36.41 $\pm$ 7.00
PC-EA 67.40 $\pm$ 1.23 31784 $\pm$ 1171 16.33 $\pm$ 0.62 33.35 $\pm$ 0.75 28.99 $\pm$ 0.75 21.33 $\pm$ 1.21
LC-100 66.49 $\pm$ 4.12 32270 $\pm$ 8551 16.35 $\pm$ 1.89 29.77 $\pm$ 3.30 26.04 $\pm$ 2.15 27.85 $\pm$ 5.37
1.0 NEA 59.23 $\pm$ 0.86 45382 $\pm$ 2627 20.99 $\pm$ 0.91 34.48 $\pm$ 2.48 26.71 $\pm$ 0.52 17.82 $\pm$ 3.30
PC-EA 57.42 $\pm$ 0.41 49016 $\pm$ 822 22.89 $\pm$ 0.60 39.39 $\pm$ 0.75 26.33 $\pm$ 0.72 11.38 $\pm$ 0.93
Table 6: Performance of networks from LC-100, NEA, and PC-EA at three $\alpha$ values, as well as the STL network. Bolded values are best for the corresponding $\alpha$. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satsified with 3 or more transfers. All values are averaged over ten random seeds, with one standard deviation following “$\pm$”, except for those in the STL row.
Figure 7: Average trip time $C_{p}$ versus total route time $C_{o}$ achieved by LC-100, NEA, and PC-EA, for $\alpha=0.0$ (bottom-right), 0.5 (middle), and 1.0 (top-left), as well as for the STL network itself (independent of $\alpha$). Except for STL, each point is a mean value over 10 random seeds, and bars show one standard deviation. Note that PC-EA’s $\alpha=0.0$ and 0.5 have nearly identical values (as shown in Table 6), so the corresponding points are stacked in this figure and appear as one point.
@@
Comparing NEA’s results to LC-100, we see that at $\alpha=0.0$, NEA decreases $C_{o}$ and increases $C_{p}$ relative to LC-100; and for $\alpha=1.0$, the reverse is true. In both cases, NEA improves over the initial network on the objective being optimized, at the cost of the other, as we would expect. In the balanced case ($\alpha=0.5$), we see that NEA decreases $C_{o}$ by 13% while increasing $C_{p}$ by 4%. This matches what we observed in Figure 3, as discussed in subsection 5.1.
@@
Looking at the transfer percentages $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$, we see that as $\alpha$ increases, the overall number of transfers shrinks, with $d_{1}$ growing and $d_{un}$ shrinking. This is as expected, since fewer transfers usually means shorter passenger travel times. By the same token, PC-EA performs best on these metrics, as it does on average trip time $C_{p}$. Notably though, at $\alpha=0.0$ NEA’s $d_{0}$, $d_{1}$, and $d_{2}$ are only slightly worse than the STL network’s - the largest increase is in $d_{2}$, by 13% of STL’s value - while $d_{un}$ is less than the STL network’s, meaning that NEA’s output networks let more passengers reach their destination in 2 or fewer transfers than the STL network. In fact, the very small decrease in $d_{0}$ implies that the larger increases in $d_{1}$ and $d_{2}$ mostly come from trips that were in $d_{un}$ in the STL network; NEA has thus decreased the average number of transfers. This is a significant improvement, given that at $\alpha=0.0$ the algorithm is unconcerned with minimizing passenger travel time.
@@
At $\alpha=0.5$, NEA’s $d_{un}$ decreases further and $d_{0}$, $d_{1}$, and $d_{2}$ correspondingly increase further. Then from $\alpha=0.5$ to $1.0$, NEA’s $d_{un}$ drops by more than half, and its $d_{0}$, $d_{1}$, and $d_{2}$ are all markedly increased, especially $d_{0}$ which is 31% greater than STL’s $d_{0}$. NEA improves on the STL network by a considerable margin at all three $\alpha$ values, even though PC-EA outperforms NEA at $\alpha=1.0$.
@@
The likely aim of Laval’s transit agency is to reduce costs while improving service quality, or at least not harming it too much. The “balanced” case, where $\alpha=0.5$, is most relevant to that aim. In this balanced case, we find that NEA proposes networks that reduce total route time by 13% versus the existing STL transit network. Some simple calculations can help understand the savings this represents. The approximate cost of operating one of Laval’s buses is 200 Canadian dollars (CAD) per hour. Suppose that each route has the same headway (time between bus arrivals) $H$. Then, the number of buses $N_{r}$ required by route $r$ is the ceiling of the route’s driving time $\tau_{r}$ divided by the headway, $\left\lceil\frac{\tau_{r}}{H}\right\rceil$. The total number of buses $N_{\mathcal{R}}$ required for a transit network $\mathcal{R}$ is then:
$N_{\mathcal{R}}=\sum_{r\in\mathcal{R}}N_{r}\approx\frac{1}{H}\sum_{r\in\mathcal{R}}\tau_{r}=\frac{C_{o}}{H}$
@@
The total cost of the system is 200 CAD×$\times N_{\mathcal{R}}$. Assuming a headway of 15 minutes on all routes, that gives a per-hour operating cost of 319,387 CAD for the STL network, while the average operating cost of networks from NEA with $\alpha=0.5$ is 279,707 CAD. So, under these simplifying assumptions, NEA could save the Laval transit agency on the order of 40,000 CAD per hour - savings of more than 12% of costs. This is only a rough estimate: in practice, headways differ between routes, and there may be additional costs involved in altering the routes from the current system, such as from moving or building bus stops. But it shows that NEA, and neural heuristics more broadly, may offer practical savings to transit agencies.